# A Probabilistic Parking Process and Labeled IDLA

### v1

@@ Abstract.
We introduce and study a new probabilistic variant of the classical parking protocol of Konheim and Weiss [29], which is closely related to Internal Diffusion Limited Aggregation, or IDLA, introduced in 1991 by Diaconis and Fulton [15]. In particular, we show that if one runs our parking protocol starting with a parking function whose outcome permutation (in the sense of the classical parking process of Konheim and Weiss) is the identity permutation, then we can compute the exact probability that all of the cars park. Furthermore, we compute the expected time it takes for the protocol to complete assuming all of the cars park, and prove that the parking process is negatively correlated. We also study statistics of uniformly random weakly increasing parking functions, a subset of parking functions whose outcome is the identity permutation. We give the distribution of the last entry, along with the probability that a specific set of cars is lucky, and the expected number of lucky cars.
@@ 1. Introduction
In 1966 Konheim and Weiss [29] introduced a now classical parking problem, which is defined as follows. Fix $n\in\mathbb{N}=\{1,2,3,\ldots\}$ and consider a sequence of $n$ cars, each given labels in $[n]\coloneqq\{1,\dots,n\}$, attempting to park on the sites of the line segment $[n]$ according to the following protocol. Initially, each car records its preferred parking spot in the preference list $\alpha\coloneqq(\alpha_{1},\dots,\alpha_{n})\in[n]^{n}$ with $\alpha_{i}$ being the preference of car $i$. The cars then enter the street from left to right, and try to park in increasing order of their labels. If their preferred parking spot is unoccupied, they are able to immediately park in their desired spot. Otherwise, each car continues to drive to the next available spot to the right and parks there. If there is no such parking spot available, the car exits the street and is unable to park. The set of preference lists $\alpha$ such that all cars can park under the aforementioned protocol are called parking functions, and the set of all parking functions with $n$ cars is denoted by $\operatorname{PF}_{n}$. To give a few simple, but illustrative examples, the preference list $(1,1,1\dots,1)$ is a parking function, as is any permutation in $S_{n}$. However, the preference list $(n,n,n\dots,n)$ is not a parking function, as the second car fails to park. Konheim and Weiss, interested in studying hash functions, initiated the study of these combinatorial objects by proving that $|\operatorname{PF}_{n}|=(n+1)^{n-1}$, and since, classical parking functions and several new variants [6, 7, 20] have been of great interest to the enumerative combinatorics community.
@@
In addition, parking functions, along with their subsets and generalizations, have connections to a vast range of mathematical fields. Past work connects parking functions to computing volumes of polytopes, hyperplane arrangements, representation theory, the QuickSort algorithm, enumerating Boolean intervals in the weak Bruhat order of the symmetric group, facet structure of the permutahedron, and even to the continued fraction expansion of $\sqrt{2}$ and to the Tower of Hanoi game [1, 17, 26, 35, 22, 20, 19, 41, 28, 24]. For a brief introduction to parking functions and some of their properties, we recommend [33]. For a detailed overview of known results about parking functions, we point the reader to Yan [44] and for many open problems involving parking functions we recommend [4].
@@
In this work, we are interested in studying probabilistic parking protocols for specific subsets of parking functions having particular outcome permutations, where the outcome permutation is the final order of the parked cars at the end of the classical parking process. In particular, we study sets of parking functions whose outcome is the identity permutation, a set which we denote $\mathrm{PF}_{n}(\mathrm{id})$. For example, $(1,2,1)\in\operatorname{PF}_{3}(\mathrm{id})$, but $(2,1,3)\notin\operatorname{PF}_{3}(\mathrm{id})$ since the cars do not park in order the identity permutation.
Note that, one well known subset of parking functions whose outcome is the identity permutation are called weakly increasing parking functions, which are known Catalan objects (i.e., combinatorial objects enumerated by the Catalan numbers). A weakly increasing parking function is exactly as it sounds: a parking function whose entries are weakly increasing. However, the following equivalent characterization of weakly increasing parking functions is more useful in our work.
@@
A weakly increasing parking function of length $n$ is a preference list $\alpha\in[n]^{n}$ which is weakly increasing, and has the property that $\alpha_{i}\leq i,$ for all cars $i\in[n]$. We denote the set of all weakly increasing parking functions by $\mathsf{PF}^{\uparrow}_{n}$.
Probabilists have steadily increased their interest in parking processes and parking functions. Recently there have been several works studying phase transitions in the parking process on random graphs such as random trees [2, 8, 9, 10, 12, 13, 23], and Frozen Erős Rényi graphs [11]. Diaconis and Hicks [16], on the other hand, considered the question “What does a random parking function look like?”. They gave probabilistic interpretations of some known generating functions, and used these results to derive statistics about parking functions.
However, there are very few works which introduce randomness into the parking protocol itself. Durmić, Han, Harris, Ribeiro, and Yin [18] give one such model. With probability $q=1-p$ the car reverses the classical parking protocol, and searches backwards in an attempt to park in the first open spot it finds to the left. In their parking protocol, whenever a car attempting to park finds its preferred parking spot occupied, with probability $p$ the car proceeds as in the classical parking protocol, and parks in the first empty spot it finds to the right. Remarkably, the probability of obtaining a parking function is shown to be independent of the forward probability $p$. They also provided some combinatorial consequences of the probabilistic parking protocol, which gave interpretations of the triangular arrays produced by a generating polynomial related to the study of duplicial algebras and Lagrange inversion [18, Theorem 8].
Motivated by the results of Diaconis and Hicks [16] and Durmić et al. [18], we study statistics of a model of stochastic parking when it is started with an initial preference list which is a parking function whose outcome is the identity permutation. Our contributions can be broadly categorized into two themes:
@@
    Statistics from two different models of stochastic parking which we describe in detail in Section 1.1 below, and
@@
    statistics of the entries of uniformly random weakly increasing parking functions.
We show that, when considering the set of parking functions whose outcome permutation is the identity permutation, both of our models of stochastic parking are almost completely characterized by the classical gambler’s ruin problem (for an overview of the subject, see [21, 25, 31], which offer excellent expositions). Therefore, our goal with (2) is to characterize properties of a subset of these parking functions which allow us to obtain succinct expressions for (1).
@@ 1.1. Models of Stochastic Parking
We now introduce our models of stochastic parking. Let $n\in\mathbb{N}$ be the number of cars and consider a preference list $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in[n]^{n}$. Upon the arrival of car $i\in[n]$, it attempts to park in its preferred spot $\alpha_{i}\in[n]$. If spot $\alpha_{i}$ is unoccupied, the car parks there. Otherwise, with probability $0\leq p\leq 1$, it displaces one spot to the right and attempts to park in spot $\alpha_{i}+1$ so long as $\alpha_{i}+1\in[n]$. Conversely, with probability $q=1-p$, it displaces one spot to the left and attempts to park in spot $\alpha_{i}-1$ so long as $\alpha_{i}-1\in[n]$. In this way, car $i$ performs a random walk starting from its preferred spot $\alpha_{i}$, and parking in the first unoccupied spot on the line segment $[n]$ it encounters, if any. Note that the case in which $p=1$ reduces to the original protocol of Konheim and Weiss [29], while the case of $p=0$ corresponds to running the classical protocol from the right end of the segment to the left. As in the classical model, cars arrive one at a time, and a car may enter the street only after its predecessor has completed its search process. Now, we distinguish two different variants of this model depending on the dynamics at the boundaries, namely spots $1$ and $n$:
@@
    Open boundary. Cars are free to drift around the finite set $[n]$. If a car exits this set (i.e., if it enters the set $\mathbb{Z}\setminus[n]$), it immediately terminates its search process and is unable to park. In particular, if a car is in spot $n$ (resp., spot $1$) and randomly displaces to spot $n+1$ (resp., spot 00), it cannot re-enter the set $[n]$.
Therefore, the central question in this setting is not that of membership in the set $\operatorname{PF}_{n}$, but rather the probability with which a preference list $\alpha\in[n]^{n}$ allows all cars to park.
@@ 1.2. Contributions
In our work we consider two subsets of parking functions, the set of weakly increasing and a superset consisting of those parking functions with outcome the identity permutation, which can be characterized as follows.
@@
A preference list $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in[n]^{n}$ satisfies $\alpha\in\operatorname{PF}_{n}(\mathrm{id})$ if and only if its entries satisfy $\alpha_{i}\leq i$ for all $i\in[n]$.
@@
|$\left|\left\{k:\pi_{k}\leq i\right\}\right|\geq i,\quad\quad\mbox{for all }1\leq i\leq n\,.$
Thus, necessarily, if $\alpha_{i}\leq i$ for all $i\in[n]$, $\alpha$ must be a parking function of length $n$. Furthermore, if $\alpha_{i}\leq i$ for every $i\in[n]$, then we may proceed by induction. It is clear that $\alpha_{1}=1$. Next, assume that for some $1\leq j<n$, every $k$th car, where $1\leq k\leq j$ parks in spot $k$. Then since $\alpha\in\operatorname{PF}_{n}$ the $(k+1)$th car parks in some spot $k<b_{k+1}\leq k+1$, hence $b_{k+1}=k+1$ and the result holds. ∎
For both models, we give expressions for the probability of parking and expected time to park (i.e., a random variable analogue of the displacement statistic), conditioned on ultimately parking. In the following, for each $i\in[n]$, we let $d_{i}\coloneqq i-\alpha_{i}$ denote the deterministic displacement of car $i$ for a parking function $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in\mathrm{PF}_{n}(\mathrm{id})$. In addition, we let $X_{\alpha}^{i}$ be an indicator random variable which is $1$ if car $i$ parks under the stated protocol and preference list $\alpha$, and 00 otherwise. Similarly, we let $X_{\alpha}$ be an indicator random variable which is $1$ if all of the cars park under the stated protocol and preference list $\alpha$, and 00 otherwise. Let $Z_{t}^{i}$ be the random variable keeping track of the position of car $i$ at time $t$, and for each car, we define the stopping time
@@
The proofs of Theorems 1.3 and 1.4 are based on gambler’s ruin arguments. Theorem 1.3 in particular invokes results from McMullen [34]. Notice, in contrast to the model of Durumić et al., [18] the probability that all of the cars park in our case retains the parameter $p$. Moreover, in Section 5 we prove that the random variables $X_{\alpha}^{i}$ are negatively correlated.
@@
Let $\alpha\sim\mathcal{U}(\operatorname{PF}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ sufficiently large, and fixed $j\geq 1$,
(1.1) $\Pr_{\alpha\,\sim\,\mathcal{U}(\operatorname{PF}_{n})}\left[\alpha_{n}=n-j\right]\sim\frac{1}{\sqrt{2}}\frac{j-1}{4^{n}}.$
@@ 1.3. Related Work
The probabilistic parking protocol described can easily be interpreted as an interacting particle system (for a thorough treatment of the subject, refer to [32, 39]). In particular, preference lists translate directly to initial configurations of particles (i.e., cars) on the line segment such that multiple particles are allowed on the same site.
@@
The underlying Markov chain would have state space $[n]^{n}$, and the set of initial configurations are those allowing each site to contain a non-negative number of labeled particles, with the set of absorbing states being those in which each site contains zero or one particle.
@@
More specifically, one can interpret the dynamics of our probabilistic parking process as being a special case of the model known as Internal Diffusion Limited Aggregation, or IDLA, which was originally introduced by Diaconis and Fulton [15, 30]. IDLA is also known to be equivalent to a special case of the Activated Random Walk model when the rate of a particle sleeping $\lambda$ is set to be infinite [38].
These models are of interest to the interacting particle system community as they exemplify a phenomena called self-organized criticality in which the system appears to undergo a phase transition, not by way of fine tuning the parameters of the model, but rather as a natural result of the dynamics of the process.
@@
Our model differs from these more classical models, however, as we enforce that the cars, or particles, remain distinct throughout the process. That is we give a total order to the labels of the cars attempting to park, and so when several cars are assigned to the same initial spot, the car with the smallest number in the order is allowed to park immediately while other cars must perform random walks in increasing order of their labels. This seems to have the effect of making the model more sensitive to the initial configuration $\alpha$ than the usual IDLA model in one dimension.
@@
Furthermore, recently and independently Varin introduced the Golf Model [43], an interacting particle system that can be seen as a generalization of IDLA. In their work, Varin also introduces a probabilistic parking protocol on the cycle $\mathbb{Z}/n\mathbb{Z}$ as a special case of the Golf Model in which each car starts at a uniformly random spot on the circular parking lot. The cars move to the right with probability $p$ and to the left with probability $1-p$ to find the nearest open parking spot, as in our process. However, the work of Varin differs in several ways. Perhaps most significantly, they focus on the setting in which there are strictly more parking spots than cars. Thus, they are interested in studying the distribution of empty parking spots remaining at the end of the process. Furthermore, the parking process of Varin is essentially equivalent to the IDLA model, and as such, lets the cars be indistinguishable.

### v2

@@ Abstract.
In 1966, Konheim and Weiss [konheimWeiss] introduced a now classical parking protocol. The deterministic process and its resultant objects, known as parking functions, have since become a favorite object of study in enumerative combinatorics. In our work, we introduce and study a probabilistic variant of the classical parking protocol, which is closely related to Internal Diffusion Limited Aggregation, or IDLA, introduced in 1991 by Diaconis and Fulton [diaconis1991growth]. In particular, we compute the stationary distribution of this process when initiated with a particular class of initial preferences, of which weakly increasing parking functions are a subset. Furthermore, we compute the expected time it takes for the protocol to complete assuming all of the cars park, and prove that, in some cases, the parking process is negatively correlated. In addition, we study statistics of uniformly random weakly increasing parking functions such as the distribution of the last entry, the probability that a specific set of cars is lucky, and the expected number of lucky cars.
@@ 1. Introduction
In 1966 Konheim and Weiss [konheimWeiss] introduced a now classical parking problem which is defined as follows. Fix $n\in\mathbb{N}=\{1,2,3,\ldots\}$ and consider a sequence of $n$ cars, each given labels in $n]\coloneqq\{1,2,\dots,n\}$, attempting to park on the sites of the line segment $[n]$ according to the following protocol. Initially, each car records its preferred parking spot in the preference list $\alpha\coloneqq(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in[n]^{n}$, with $\alpha_{i}$ being the preference of car $i$. The cars then enter the street from left to right, and attempt to park in increasing order of their labels. If their preferred parking spot is unoccupied, they are able to immediately park in their desired spot. Otherwise, each car continues to drive to the next available spot to the right and parks there. If there is no such parking spot available, the car exits the street and is unable to park. The set of preference lists $\alpha$ such that all cars can park under the aforementioned protocol are called parking functions, and the set of all parking functions with $n$ cars is denoted by $\mathsf{PF}_{n}$. To give a few simple but illustrative examples, the preference list $(1,1,1,\dots,1)$ is a parking function, as is any permutation of $[n]$. We let $S_{n}$ denote the set of permutations of $[n]$, and we write permutations in one-line notation. However, the preference list $(n,n,n,\dots,n)$ is not a parking function, as the second car already fails to park. Konheim and Weiss, interested in studying hash functions, initiated the study of these combinatorial objects by proving that $|\mathsf{PF}_{n}|=(n+1)^{n-1}$. Since then, classical parking functions and several new variants [colaric2020interval, countingKnaples, fang2024vacillatingparkingfunctions] have been of great interest to the enumerative combinatorics community.
@@
In addition, parking functions, along with their restrictions and generalizations, have connections to a vast range of mathematical fields. Past work connects parking functions to computing volumes of polytopes, hyperplane arrangements, representation theory, sorting algorithms, enumerating Boolean intervals in the weak Bruhat order of the symmetric group, facet structure of the permutahedron, and even to the continued fraction expansion of $\sqrt{2}$ and to the Tower of Hanoi game [Aguillon2022OnPF, MR3724106, Harris2023LuckyCA, unit_perm, garcia2024defectiveparkingfunctionsyoung, fang2024vacillatingparkingfunctions, elder2025parking, MR1902680, HicksDiagonalHarmonics, MR3551576]. For a brief introduction to parking functions and some of their properties, we recommend [mori2024parkingfunction]. For a detailed overview of known results about parking functions, we point the reader to Yan [yan2015parking], and for many open problems involving parking functions we recommend [carlson2020parking].
@@
In this work, we will be interested in discussing the behavior of a probabilistic parking protocol when initiated with a preference vector which is a parking function (in the classical sense) having a specific outcome permutation, where the outcome permutation is the final order of the parked cars at the end of the classical parking process. In particular, we will consider parking functions whose outcome is the identity permutation, a set which we denote $\mathrm{PF}_{n}(\mathrm{id})$. For example, $(1,2,1)\in\mathsf{PF}_{3}(\mathrm{id})$, but $(2,1,3)\notin\mathsf{PF}_{3}(\mathrm{id})$ since the cars do not park in the order of the identity permutation.
We will also discuss statistics of weakly increasing parking functions, which are parking functions whose entries are weakly increasing. More formally, we have the following definition.
@@
A weakly increasing parking function of length $n\geq 1$ is a preference list $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in[n]^{n}$ such that $\alpha_{i}\leq i$ for all $i\in[n]$ and $\alpha_{1}\leq\alpha_{2}\leq\cdots\leq\alpha_{n}$. We denote the set of weakly increasing parking functions of length $n\geq 1$ by $\mathsf{PF}^{\uparrow}_{n}$.
Weakly increasing parking functions are known Catalan objects (i.e., combinatorial objects enumerated by the Catalan numbers, for more on Catalan objects see [catalan_stanley]). Moreover, the outcome of any weakly increasing parking function is the identity permutation (i.e., $\mathsf{PF}^{\uparrow}_{n}\subseteq\mathrm{PF}_{n}(\mathrm{id})$). However, note that the converse is not true: for example, $(1,2,1)\in\mathsf{PF}_{3}$ parks in the order $123$, but is not weakly increasing.
The classical parking process, and resultant parking functions, have steadily increased in popularity as objects of study in the probability community. Recently there have been several works studying phase transitions in the parking process on random graphs such as random trees [aldous2023parking, contat2022sharpness, contat2023parking, contat2023parking1, contat2025parking, curien2019phase, goldschmidt2019parking], and Frozen Erős-Rényi graphs [ContatCurien]. Diaconis and Hicks [diaconis2017probabilizing], on the other hand, considered the question “What does a random parking function look like?”. They gave probabilistic interpretations of some known generating functions, and used these results to derive statistics about parking functions.
However, there are relatively few works which introduce randomness into the parking protocol itself. Durmić, Han, Harris, Ribeiro, and Yin [durmic2022probabilistic] give one such model. In their parking protocol, whenever a car attempting to park finds its preferred parking spot occupied, with probability $p$ the car proceeds as in the classical parking protocol, and parks in the first empty spot it finds to the right. With probability $q\coloneqq 1-p$ the car reverses the classical parking protocol, and searches backwards in an attempt to park in the first open spot it finds to the left. Remarkably, the probability of obtaining a parking function is shown to be independent of the forward probability $p$. They also provided some combinatorial consequences of the probabilistic parking protocol, which gave interpretations of the triangular arrays produced by a generating polynomial related to the study of duplicial algebras and Lagrange inversion [durmic2022probabilistic, Theorem 8].
Motivated by the results of Diaconis and Hicks [diaconis2017probabilizing] and Durmić et al. [durmic2022probabilistic], we study statistics of a model of stochastic parking when it is started with an initial preference list which is a parking function whose outcome is the identity permutation. Note, that there are $n!$ parking functions with outcome the identity permutation, namely those preferences $\alpha\coloneqq(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$ satisfying $a_{i}\leq i$ for all $i\in[n]$. For a proof that this property indeed characterizes the elements of $\mathrm{PF}_{n}(\mathrm{id})$, see Lemma 1.3 below. Our contributions can be broadly categorized into two themes:
@@
    Analyze properties of two different models of stochastic parking which we describe in detail in Section 1.1 below, and
@@
    study statistics of the entries of uniformly random weakly increasing parking functions.
We show that, when considering the set of parking functions whose outcome permutation is the identity permutation, both of our models of stochastic parking are almost completely characterized by the classical gambler’s ruin problem (for an overview of the subject, see [feller1968introduction, grinstead2006grinstead, levinPeresWilmer], which offer excellent expositions). Therefore, our goal with (2) is to characterize properties of a subset of these parking functions which allow us to obtain succinct expressions for (1).
@@ 1.1. Models of Stochastic Parking
We now introduce our models of stochastic parking. Let $n\in\mathbb{N}$ be the number of cars and consider a preference list $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in[n]^{n}$. As in the classical model, cars arrive one at a time, and a car may enter the street only after its predecessor has completed its search process. Upon the arrival of car $i\in[n]$, it attempts to park in its preferred spot $\alpha_{i}\in[n]$. If spot $\alpha_{i}$ is unoccupied, the car parks there. Otherwise, with probability $p\in[0,1]$, it displaces one spot to the right and attempts to park in spot $\alpha_{i}+1$ so long as $\alpha_{i}+1\in[n]$. Conversely, with probability $q\coloneqq 1-p$, it displaces one spot to the left and attempts to park in spot $\alpha_{i}-1$ so long as $\alpha_{i}-1\in[n]$. In this way, car $i$ performs a random walk starting from its preferred spot $\alpha_{i}$, and parking in the first unoccupied spot on the line segment $[n]$ it encounters, if any. Note that the case in which $p=1$ reduces to the original protocol of Konheim and Weiss [konheimWeiss], while the case of $p=0$ corresponds to running the classical protocol from the right end of the segment to the left. Now, we distinguish two different variants of this model depending on the dynamics at the boundaries, namely spots $1$ and $n$:
@@
    Open boundary. Cars are free to drift around the finite set $[n]$. If a car exits this set (i.e., if it enters the set $\mathbb{Z}\setminus[n]$), it immediately terminates its search process and is unable to park. In particular, if a car is in spot $n$ (resp., spot $1$) and randomly displaces to spot $n+1$ (resp., spot 0), it cannot re-enter the set $[n]$.
Our question of interest in this context is the following.
Question 1.2.
Fix $\alpha\in[n]^{n}$ and $\sigma\in S_{n}$. What is the probability that the probabilistic parking protocol iniated with preferences $\alpha$ allows all of the cars to park with outcome $\sigma$?
[…]
Figure 1. Illustration of the probabilistic parking protocol as a random mapping from $n]^{n}$ to $S_{n}$. The dynamics induce a distribution over outcome permutations, and our main question of interest is to characterize this distribution $\Pr\left[\operatorname{Out}(\alpha)=\sigma\right]$ where $\alpha$ and $\sigma$ are fixed, and we use $\operatorname{Out}(\alpha)$ to denote the outcome of $\alpha$ under the probabilistic parking protocol.
@@ 1.2. Contributions
In our work we consider two subsets of parking functions, the set of weakly increasing parking functions and a superset consisting of those parking functions with outcome the identity permutation, which can be characterized as follows.
@@
A preference vector $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in[n]^{n}$ satisfies $\alpha\in\mathsf{PF}_{n}(\mathrm{id})$ if and only if its entries satisfy $\alpha_{i}\leq i$ for all $i\in[n]$.
@@
$\left|\left\{k:\alpha_{k}\leq i\right\}\right|\geq i,\quad\quad\mbox{for all }1\leq i\leq n\,.$
Thus, necessarily, if $\alpha_{i}\leq i$ for all $i\in[n]$, $\alpha$ must be a parking function of length $n$. Furthermore, if $\alpha_{i}\leq i$ for every $i\in[n]$, then we may proceed by strong induction. It is clear that $\alpha_{1}=1$. Next, assume that for some $1\leq j<n$, every $k$th car, where $1\leq k\leq j$ parks in spot $k$. Then since $\alpha\in\mathsf{PF}_{n}$ the $(k+1)$th car parks in some spot $k<b_{k+1}\leq k+1$, hence $b_{k+1}=k+1$ and the result holds. ∎
Notice that, not only does this characterize the set of parking functions with outcome permutation being the identity in the classical parking protocol of Konheim and Weiss, but a similar proof shows that the same condition characterizes the set of preferences in which, if all of cars are to park in the probabilistic parking protocol (with either boundary condition), then they must do so in the order of the identity permutation. In other words, there are no other (and no fewer) preferences which would allow all of the cars to park in the identity configuration in the probabilistic setting defined above. See Figure  1 for an illustration of the idea. For both of our probabilistic models, we give expressions for the probability of parking and expected time to park (i.e., a random variable analogue of the displacement statistic), assuming we start with a preference vector $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$ and conditioned on ultimately parking. Namely, we answer Question 1.2 for two subsets of preferences, those which are parking functions with outcome being the identity permutation, and those with outcome being the reverse of the identity.
In the following, for each $i\in[n]$, we let $d_{i}\coloneqq i-\alpha_{i}$ denote the deterministic displacement of car $i$ for a parking function $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in\mathrm{PF}_{n}(\mathrm{id})$. In addition, we let $X_{\alpha}^{i}$ be an indicator random variable which is $1$ if car $i$ parks under the stated protocol and preference list $\alpha$, and 0 otherwise. Similarly, we let $X_{\alpha}$ be an indicator random variable which is $1$ if all of the cars park under the stated protocol and preference list $\alpha$, and 0 otherwise. Let $Z_{t}^{i}$ be the random variable keeping track of the position of car $i$ at time $t$, and for each car, we define the stopping time
@@
The proofs of Theorems 1.4 and 1.5 are based on gambler’s ruin arguments. Theorem 1.4 in particular invokes results from McMullen [mcmullen2020drunken]. Notice, in contrast to the model of Durmić et al., [durmic2022probabilistic] the probability that all of the cars park in our case retains the parameter $p$. Moreover, in Section 5 we prove that the random variables $X_{\alpha}^{i}$ are negatively correlated.
@@
Let $\alpha\sim\mathcal{U}(\mathsf{PF}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ sufficiently large, and fixed $j\geq 1$,
(1.1) $\Pr_{\alpha\,\sim\,\mathcal{U}(\mathsf{PF}_{n})}\left[\alpha_{n}=n-j\right]\sim\frac{1}{\sqrt{2}}\frac{j-1}{4^{n}}.$
@@ 1.3. Related Work
The probabilistic parking protocol described can easily be interpreted as an interacting particle system (for a thorough treatment of the subject, refer to [Liggett_2004, SPITZER1970246]). In particular, preference vectors translate directly to initial configurations of particles (i.e., cars) on the line segment such that multiple particles are allowed on the same site.
@@
The underlying Markov chain would have state space $n]^{n}$, and the set of initial configurations are those allowing each site to contain a non-negative number of labeled particles, with the set of absorbing states being those in which each site contains zero or one particle.
@@
More specifically, one can interpret the dynamics of our probabilistic parking process as being a generalization of the model known as Internal Diffusion Limited Aggregation, or IDLA, which was originally introduced by Diaconis and Fulton [diaconis1991growth, lawler]. (IDLA is also known to be essentially equivalent to a special case of the Activated Random Walk model when the rate of a particle sleeping $\lambda$ is set to be infinite [rolla2020activated].)These models are of interest to the interacting particle system community as they exemplify a phenomena called self-organized criticality in which the system appears to undergo a phase transition, not by way of fine tuning the parameters of the model, but rather as a natural result of the dynamics of the process.
@@
Our model differs from this more classical model, however, as we enforce that the cars (or particles) remain distinct throughout the process, and we are interested in the positions of the labeled cars at the end of the process. That is we give a total order to the labels of the cars attempting to park, and so when several cars are assigned to the same initial spot, the car with the smallest number in the order is allowed to park immediately while other cars must perform random walks in increasing order of their labels. One can then see our result then as working towards characterizing the stationary distribution of the Markov chain. In general, computing stationary measures for various interacting particle systems is a classical, and still lively, area of research at the intersection of probability theory and combinatorics (e.g., [bramson2002characterization, CorteelSylvie2011Tcft, MartinStationary, Nestoridi_Schmid_2024, yang2023stationarymeasureshigherspin]).
@@
Recently and independently Varin introduced the Golf Model [VarinGolf], an interacting particle system that can be seen as a generalization of IDLA. In their work, Varin also introduces a probabilistic parking protocol on the cycle $\mathbb{Z}/n\mathbb{Z}$ as a special case of the Golf Model in which each of the, now unlabeled, cars starts at a uniformly random spot on the circular parking lot. The cars move to the right with probability $p$ and to the left with probability 1p1-p to find the nearest open parking spot, as in our process. The focus of our works differ, however, as Varin is interested in the setting in which there are strictly more parking spots than cars. Thus, they are interested in studying the distribution of the set of unoccupied sites at the end of the process.
In publishing this work, we were made aware that Nadeau and Tewari also investigate a general form of IDLA on $\mathbb{Z}$, which is essentially equivalent to our model with open boundaries. In [nadeau2021qdeformationalgebraklyachkomacdonalds] the authors study at a generalization of the classical IDLA model with general jump rates, and determine the probability that the final configuration of the unlabeled particles on the line is to occupy the set of sites in $[n]$. The authors approach is strictly algebraic, and they are instead interested in enumerative results related to this probability, obtaining an expression for it in terms of remixed Eulerian numbers [NadeauRemixedEulerian]. More recently (and independently) Nadeau formalized this interacting particle as a parking procedure in [nadeau:hal-04262134].
Overall, one can view our contribution to this literature as investigating the model from a probabilistic perspective and obtaining explicit expressions for the probability of parking in terms of the parameters $p$ and $\alpha$, while still maintaining an understanding of where the cars end up parking. This allows us to adopt the perspective of the parking process as a random mapping from $n]^{n}$ to $S_{n}$. Moreover we offer new computations for the expected time to park, as well as introduce new boundary conditions for the process.